# ClassInvGen: Class Invariant Synthesis using Large Language Models

### v1

@@ 4.1. Benchmark
We use a C++ implementation of classical data structures for our micro-benchmarks (Dimitrios, 2023), which include 9 data structures with associated unit tests: AvlTree, BinarySearchTree, HashTable, Heap, LinkedList, Queue, RedBlackTree, Stack, and Vector. Table 1 shows the statistics of these benchmark data structures. To ensure correctness, we thoroughly examined each benchmark example and corrected a few implementation bugs, treating this refined benchmark as the ground truth. All subsequent experiments are based on this benchmark setup.
In addition to textbook examples, we also conducted experiments on utility classes (z3Prover, 2024) from Z3 (de Moura and Bjørner, 2008), including ema, dlist, heap, hashtable, permutation, scoped_vector, and the most complex class, bdd_manager. The latter will be discussed in detail as a case study, including an evaluation with one of the authors of Z3.
4.2. Evaluation
@@
  1. (1)
    How many of these invariants are correct (with respect to the user-provided test cases) and do they capture essential properties of the source code (Section 4.3)?
  2. (2)
@@
Figure 12. Evaluation of ClassInvGen generated invariants
4.3. Correctness
ClassInvGen produces filtered invariants, which we evaluate using an automated pipeline (Figure 12) to assess their quality against the data structures benchmark described in Section 4.1. The evaluation process begins by instrumenting the source code with the candidate invariants, as detailed in Section 3.1.3. Unlike the generated tests used in ClassInvGen for filtering, the evaluation relies on unit tests containing input-output pairs. A filtered invariant is considered correct if it does not report any errors for any tests that successfully compile and run. Notably, these unit tests are not available during the ClassInvGen generation phase and are used exclusively for evaluation purposes.
For each passing filtered invariant, the authors of the paper conducted a manual review. This examination confirmed that all validated filtered invariants are indeed correct and effectively capture properties of the data structures. e.g., all 3 properties of AvlTree described in Section 2 are found in ClassInvGen invariants.
@@ 4.3. Correctness
# invariants # compiles # pass unit tests unit test pass rate
@@
Example # filtering tests # filtered inv. # good inv. unit tests pass rate
@@
Table 3. For each example, the table shows the number and line coverage of filtering tests, the number of filtered invariants without refinement, and with 1 refinement, as well as the unit test pass rate after 1 refinement. With filtering tests and 1 refinement, ClassInvGen achieves a perfect unit test pass rate.
4.4. Completeness
Table 4. ClassInvGen Performance Over Baseline for Previously Survived Mutants. The table shows additional mutants killed by ClassInvGen compared to the baseline and percentage improvement.
Data Structure Unsolved Baseline (#(%)) Add. Solved by ClassInvGen (#) Improv. (%)
@@ 4.4. Completeness
To evaluate the completeness of ClassInvGen invariants, we use a variation of mutation testing used in prior work to evaluate the strength of correct specifications (Endres et al., 2024). Mutation testing creates copies of the code with small changes. Individual copies are called mutants. If a mutant fails tests, the mutant is said to be killed, otherwise it survives. We expect that many mutations will corrupt data structures, so the number of mutants killed by ClassInvGen invariants is an indication of the strength of the invariant, or how complete those invariants are. Note that we only consider invariants that are correct.
We use the freeware tool mutate_cpp (Lohmann and other contributors, 2017) to perform program mutation, producing between 236 and 682 mutants per program. There are three possible outcomes for each mutant: it fails to compile, it crashes during execution, or it compiles and runs successfully without errors. We focus exclusively on mutants that either compile successfully but crash during execution or survive execution without error, as these are distinguishable cases. Mutants that fail to compile remain consistent across all configurations and therefore do not contribute to meaningful differentiation.
Figure 13. Completeness Experiment Result. The 3 bars from left to right are Tests, ClassInvGen, Tests+ClassInvGen. Tests+ClassInvGen kills the most mutants.
We conducted experiments to evaluate the completeness of three configurations: unit tests only, ClassInvGen only, unit tests with ClassInvGen. The last configuration has the strongest test oracles. For ClassInvGen only, we removed all assertions checking the expected output in the unit tests, leaving only the sequence of method calls consistent with the original tests.
@@
Figure 13 indicates tests with ClassInvGen kill the most mutants, showing that instrumenting unit tests with ClassInvGen invariants enhances their ability to identify buggy implementations. Figure 14 shows HashTable a mutant that survived tests but killed by ClassInvGen invariant assert(this->_size > 0 || this->table.empty()). Figure 15 shows another example killed by ClassInvGen invariant assert(load_factor > 0).
@@
Figure 14. Mutant that survived unit tests but killed by ClassInvGen
@@
Figure 15. Another Mutant that survived unit tests but killed by ClassInvGen
However, mutation testing has its limitations. For example, a mutant may be equivalent to the original program (e.g., n = 0 and n &= 0), resulting in no observable behavioral difference. Additionally, certain mutants may fall outside the scope of the invariants to detect, such as when only observer methods are altered, which do not affect the program’s state directly.
4.5. Comparison of ClassInvGen v.s. Daikon
We compare ClassInvGen with Daikon, the most widely adopted dynamic invariant generator (Ernst et al., 2007).
We replaced the invariant synthesis component in Figure 3 with Daikon, using filtering tests to generate program traces. We then ran Daikon’s invariant detector on these traces with a default confidence limit equal to 0.99 (larger values yield stronger filtering). For the most comprehensive set of invariant candidates, we combined all invariants Daikon generated at each function’s ENTER and EXIT points and used ClassInvGen invariant instrumentation to identify filtered invariants using unit tests to filter. The final output is a set of Daikon invariants that successfully pass unit test execution. On average, each benchmark example has around 5 Daikon invariants; however, in some cases, 0 to 2 of these invariants may be incorrect as shown in Table 6 examined by humans.
It is important to note that Daikon invariants are not always correct, even with respect to the tests used to infer them. Daikon outputs an invariant if its confidence level exceeds a specified configuration threshold, regardless of whether the invariant holds universally. Through the manual review, we identified 7 incorrect class invariants, as summarized in Table 5. These invariants pass unit test validation because both the filtering tests and the unit tests coincidentally constructed data structures with similar parameters , producing similar program traces. Consequently, validation was unsuccessful because these Daikon invariants based on filtering test traces also happened to hold true on the unit tests.
For instance, when Vector objects are constructed in both filtering tests and ground truth unit tests, elements are consistently push_backed in increasing order. However, users of Vector are not restricted to this behavior and could, in practice, choose to push_back larger elements first, breaking this assumed invariant. Therefore, this->data[] sorted by < is not a true class invariant.
Table 5. Incorrect Daikon Invariants that pass unit tests validation.
Data Structure Invariant Description
HashTable this->_size one of {5, 10, 40}
Heap this->data._Vector_base< int, std::allocator<int> >. _M_impl._Vector_impl_data._M_start[] elements >= 1
Vector this->data[] sorted by <
Queue this->maxSize == 10
LinkedList this->tail[].data elements >= 1
Stack this->maxSize one of {10, 160, 1280}
this->maxSize == 10
Table 6. Daikon Incorrect Invariants per Benchmark
@@
Figure 16. Daikon vs. ClassInvGen Kills
Among the 40 correct class invariants generated by Daikon and reviewed by us, the most common invariants indicate that the class pointer is not null and remains unchanged throughout the object’s lifetime (27 invariants). The second most frequent type asserts that the number of elements (typically this->n) is non-negative (6 invariants). These two categories provide minimal utility in distinguishing and eliminating mutants (incorrect programs).
The most valuable invariants are those that capture meaningful constraints, such as this->n < this->maxSize in Stack and this->head < this->maxSize && this->tail < this->maxSize in Queue, as these invariants have a noticeable impact on identifying and eliminating mutants as shown in Figure 16.
This result suggests a key weakness of Daikon: it cannot differentiate between invariants that are universally true and those that hold only in the context of specific tests. We also postulate that Daikon is less mature for C++ codebases compared to Java, explaining the lack of high-quality class invariants on our benchmark. Moreover, our findings indicate that LLMs are better at capturing the concept of true ”global” or ”class” invariants, reflecting properties that are inherent to the data structure rather than incidental to the tests used.
5. Case study: Z3 bdd_manager class
Table 7. Statistics of the studied data structures in Z3
@@ 5. Case study: Z3 bdd_manager class
As a real-world case study, we apply ClassInvGen to synthesize invariants for 7 core data structures from Z3 (de Moura and Bjørner, 2008), ranging from the simple 57-line ema class to the complex 1635-line bdd_manager. The complete set includes dlist, heap, hashtable, permutation, and scoped_vector, with varying implementation complexity and the number of dependent classes as shown in Table 7. Our results were validated by one of the Z3 authors, who confirmed at least one correct and useful invariant for each studied class, with the bdd_manager class yielding 11 valuable invariants including the 2 already present from 8 generations.
@@
Z3 is a widely adopted SMT solver used in a variety of high-stakes applications requiring rigorous correctness, such as formal verification, program analysis, and automated reasoning. It is integrated into tools like LLVM (Lattner and Adve, 2004), KLEE (Cadar et al., 2008), Dafny (Leino, 2010) and Frama-C (Kirchner et al., 2015). We selected the Z3 codebase due to its stringent correctness requirements; as an SMT solver, Z3 is employed in applications demanding high reliability. This high-stakes environment makes Z3 an ideal testbed for assessing the effectiveness of synthesized invariants.
@@
The bdd_manager class [^2] https://github.com/Z3Prover/z3/blob/master/src/math/dd/dd_bdd.h is particularly noteworthy. It was chosen because it is a self-contained example with developer-written unit tests for validation, presenting a realistic yet manageable challenge. Note that the existing developer tests were used after invariants were generated, not as input to the LLM. The bdd_manager class in Z3 is a utility for managing Binary Decision Diagrams (BDDs), which are data structures used to represent Boolean functions efficiently. In BDDs, Boolean functions are represented as directed acyclic graphs, where each non-terminal node corresponds to a Boolean variable, and edges represent the truth values of these variables (true or false). This representation simplifies complex Boolean expressions and enables efficient operations on Boolean functions.
@@
With 382 lines of code in its header and 1253 lines in the implementation file, bdd_manager surpasses standard data structure complexity, offering an opportunity to evaluate ClassInvGen’s capability to generate meaningful invariants relevant to real-world scenarios. ClassInvGen achieves this by compositional generation, recursively traversing the source program’s AST (Section 3.1.2). Recursive generation became crucial when handling large classes like bdd_manager, which exceeded the LLM’s context window. Decomposing and processing its components separately allowed us to fit relevant parts into the model’s input, demonstrating the utility of recursive invariant generation for large codebases. This supports its relevance in real-world applications beyond the benchmarks.
@@
The bdd_manager class includes a developer-written member function for checking its well-formedness, as shown in Figure 17, which we removed during ClassInvGen generation. Of the 56 invariants generated by ClassInvGen, one of Z3 main authors identified 11 distinct correct and useful invariants (e.g., Figure 18) including the 2 developer-written invariants; these invariants could potentially be integrated into the codebase. An additional 5 distinct ok invariants (e.g., Figure 19) are labeled correct but have limited utility, 16 distinct correct but useless invariants (e.g., those already checked during compilation, such as type checks and constants, Figure 20), and 2 incorrect invariants (e.g., Figure 21). The remaining invariants were repetitions within these categories. This evaluation aligns with ClassInvGen’s validation results, as our validation pipeline also identified 2 incorrect invariants that failed bdd_manager unit tests.
@@ 24 UNREACHABLE();
Figure 17. Z3 developer-written class invariants for bdd_manager class
@@
Figure 18. Correct and useful invariant for bdd_manager class
@@
Figure 19. Ok invariant for bdd_manager class
@@
Figure 20. Correct and useless invariant for bdd_manager class
@@
Figure 21. Incorrect invariant for bdd_manager class
6. Related Work
@@
In this section, we discuss how ClassInvGen relates to previous works on synthesizing program invariants statically, dynamically and neurally.
6.1. Static approaches
@@
Static techniques, such as interpolation (McMillan, 2004) or abstract interpretation (Cousot and Cousot, 1977) perform a symbolic analysis of source code to compute static over-approximations of runtime behavior and represent them as program invariants over suitable domains. These techniques are often used to prove the safety properties of the code. They focus on synthesizing loop invariants and method pre/postconditions, and a few around module-level specifications (Lahiri et al., 2009). Given the undecidability of program verification, these techniques scale poorly for real-world programs, especially in the presence of complex data structures and frameworks. In contrast, ClassInvGen can be applied to large codebases to synthesize high-quality class invariants but does not guarantee soundness by construction.
6.2. Dynamic approaches
@@
Dynamic synthesis techniques, such as Daikon (Ernst et al., 2007), DIG (Nguyen et al., 2012), SLING (Le et al., 2019), and specification mining (Ammons et al., 2002), learn invariants by observing the dynamic behaviors of programs over a set of concrete execution traces. One advantage of these dynamic techniques is that they can be agnostic to the code and generally applicable to different languages. However, these approaches are limited by the templates or patterns over which the invariants can be expressed. DySy (Csallner et al., 2008) employs dynamic symbolic execution to alleviate the problem of fixed templates for bounded executions but resorts to ad-hoc abstraction for loops or recursion. Hellendoorn et al. (2019) trained models to predict the quality of invariants generated by tools such as Daikon, but do not generate new invariants. SpecFuzzer (Molina et al., 2022) generates numerous candidate assertions via fuzzing to construct templates and filters them using Daikon and mutation testing. Finally, Geminus (Boockmann and Lüttgen, 2024) aims at synthesizing sound and complete class invariants representing the set of reachable states, guiding their search using random test cases termed Random Walk.
@@
Unlike these approaches, ClassInvGen can generate a much larger class of invariants, leveraging multimodal inputs, including source code, test cases, comments, and even the naming convention learned from training data, to enhance invariant synthesis. Further, unlike prior dynamic approaches, LLM-based test generation (an active area of research (Lemieux et al., 2023; Schäfer et al., 2023; Yang et al., 2024)) reduces the need to have a high-quality test suite to obtain the invariants.
@@
For the use case of static verification, learning-based approaches have been used to iteratively improve the quality of the synthesized inductive invariants (Garg et al., 2014, 2016; Padhi et al., 2016) from dynamic traces. However, these approaches have not been evaluated in real-world programs due to the need for symbolic reasoning.
6.3. Neural approaches
@@
LLM-based invariant synthesis is an emerging area of research with some noteworthy recent contributions. Pei et al. (2023) trained a model for zero-shot invariant synthesis, which incurs high training costs and lacks feedback-driven repair. Their approach uses Daikon-generated invariants as both training data and ground truth, which can lead to spurious invariants.
@@
Prior work on nl2postcond (Endres et al., 2024) prompts LLMs to generate pre and postcondition of Python and Java benchmarks, illustrating LLMs’ ability to generate high-quality specifications. However, they do not prune incorrect invariants and do not generate class invariants that ClassInvGen does. It is an interesting future work to combine this work with ClassInvGen to generate complete class-level specifications including pre and postconditions for the public methods of the class.
@@
For static verification, recent works include the use of LLM for intent-formalization from natural language (Lahiri, 2024), and inferring specifications and inductive program invariants (Kamath et al., 2023; Ma et al., 2024). None of these techniques scale to real-world programs due to the need for complex symbolic reasoning.
7. Limitation

### v2

@@ 4.1 Benchmark
We use a C++ implementation of classical data structures for our micro-benchmarks [9], which include 9 data structures with associated unit tests: AvlTree, BinarySearchTree, HashTable, Heap, LinkedList, Queue, RedBlackTree, Stack, and Vector. Table 1 shows the statistics of these benchmark data structures. To ensure correctness, we thoroughly examined each benchmark example and corrected a few implementation bugs, treating this refined benchmark as the ground truth. All subsequent experiments are based on this benchmark setup.
In addition to textbook examples, we also conducted experiments on utility classes [46] from Z3 [34], including ema, dlist, heap, hashtable, permutation, scoped_vector, and the most complex class, bdd_manager. The latter will be discussed in detail as a case study, including an evaluation with one of the authors of Z3.
4.2 Evaluation
@@
  1. 1.
    How many of these invariants are correct (with respect to the user-provided test cases), and do they capture essential properties of the source code (Section 4.3)?
  2. 2.
@@
Figure 4: Evaluation of ClassInvGen generated invariants
4.3 Correctness
ClassInvGen produces filtered invariants, which we evaluate using an automated pipeline (Figure 4) against our benchmark (Section 4.1). A filtered invariant is considered correct if it reports no errors for any tests that successfully compile and run. Our manual review confirmed that all validated filtered invariants are indeed correct, capturing essential properties of the data structures.
@@ 4.3 Correctness
Example # inv. # compiles # pass tests pass rate
@@
Example # filter tests # filtered inv. # good inv. pass rate
@@
Table 3: For each example, the table shows the number and line coverage of filtering tests, the number of filtered invariants without refinement, and the number of invariants after 1 refinement attempt. With filtering tests and 1 refinement, ClassInvGen achieves a perfect unit test pass rate.
4.4 Completeness
Table 4: ClassInvGen The table shows the number of mutants that survive unit tests alone (Unsolved Base) as well as the number of additional mutants killed by ClassInvGen.
Data Structure Unsolved Base (#(%)) Add. by ClassInvGen (#) Impr. (%)
@@ 4.4 Completeness
To evaluate completeness, we use mutation testing. We generate mutants using mutate_cpp [29], producing between 236 and 682 mutants per program. We focus on mutants that either compile successfully but crash during execution or survive execution without errors.
Figure 5: Completeness Experiment Result. The 3 bars from left to right are Tests, ClassInvGen, Tests+ClassInvGen. Tests+ClassInvGen kills the most mutants.
@@
Figure 5 shows how many mutants are killed with each configuration for each benchmark. Combining tests with ClassInvGen kills the most mutants. Figures 6 and 7 show examples of mutants that survived unit tests but were killed by ClassInvGen invariants.
@@
Figure 6: Mutant that survived unit tests but was killed by ClassInvGen
@@
Figure 7: Another Mutant that survived unit tests but was killed by ClassInvGen
4.5 Comparison of ClassInvGen vs. Daikon
We compared ClassInvGen with Daikon [11], using filtering tests to generate program traces for Daikon’s invariant detector. On average, each benchmark example has around 5 Daikon invariants, with some being incorrect (see summary in Table 5 and complete results of invariants in Appendix 0.B).
Table 5: Daikon Incorrect Invariants per Benchmark
@@
Through manual review, we identified 7 incorrect Daikon invariants that pass unit test validation. These invariants pass because both the filtering tests and unit tests coincidentally constructed similar data structures.
Figure 8: Daikon vs. ClassInvGen Kills
Most Daikon invariants simply indicate that class pointers are not null (27 of 40 correct invariants) or that element counts are non-negative (6 invariants). The most valuable invariants, like this->n < this->maxSize in Stack, have more impact on identifying mutants (Figure 8).
This shows a key weakness of Daikon: it cannot differentiate between universally true invariants and those that hold only in specific test contexts. LLMs are better at capturing true “class” invariants that are inherent to the data structure rather than incidental to the tests.
5 Case study: Z3 bdd_manager class
Table 6: Statistics of the studied data structures in Z3
@@ 5 Case study: Z3 bdd_manager class
As a real-world case study, we apply ClassInvGen to synthesize invariants for 7 core data structures from Z3 [34], ranging from the simple 57-line ema class to the complex 1635-line bdd_manager. The complete set includes dlist, heap, hashtable, permutation, and scoped_vector, with varying implementation complexity and the number of dependent classes as shown in Table 6. Our results were validated by one of the Z3 authors, who confirmed at least one correct and useful invariant for each studied class, with the bdd_manager class yielding 11 valuable invariants including the 2 already written by Z3 authors.
@@
Z3 is a widely adopted SMT solver used in a variety of high-stakes applications that depend on its correctness. It is integrated into tools like LLVM [23], KLEE [6], Dafny [26] and Frama-C [20]. We selected the Z3 codebase due to its stringent correctness requirements; as an SMT solver, Z3 is employed in applications demanding high reliability. This high-stakes environment makes Z3 an ideal testbed for assessing the effectiveness of synthesized invariants.
@@
The bdd_manager class [^2] https://github.com/Z3Prover/z3/blob/master/src/math/dd/dd_bdd.h is particularly noteworthy. It was chosen because it is a self-contained example with developer-written unit tests for validation, presenting a realistic yet manageable challenge. Note that the existing developer tests were used after invariants were generated, not as input to the LLM. The bdd_manager class in Z3 is a utility for managing Binary Decision Diagrams (BDDs) [4], which are data structures used to represent Boolean functions efficiently. In BDDs, Boolean functions are represented as directed acyclic graphs, where each non-terminal node corresponds to a Boolean variable, and edges represent the truth values of these variables (true or false). This representation simplifies complex Boolean expressions and enables efficient operations on Boolean functions.
@@
With 382 lines of code in its header and 1253 lines in the implementation file, bdd_manager surpasses standard data structure complexity, offering an opportunity to evaluate ClassInvGen’s capability to generate meaningful invariants relevant to real-world scenarios. ClassInvGen achieves this by compositional generation, recursively traversing the source program’s AST. Recursive generation became crucial when handling large classes like bdd_manager, which exceeded the LLM’s context window. Decomposing and processing its components separately allowed us to fit relevant parts into the model’s input, demonstrating the utility of recursive invariant generation for large codebases. This supports its relevance in real-world applications beyond the benchmarks.
@@
The bdd_manager class has a developer-written member function for checking its well-formedness, as shown in Figure 9, which we removed during ClassInvGen generation. Of the 56 invariants generated by ClassInvGen, one of Z3’s main authors identified 11 distinct correct and useful invariants (e.g., Figure 10) including the 2 developer-written invariants; these invariants could potentially be integrated into the codebase. An additional 5 distinct ok invariants (e.g., Figure 11) are labeled correct but have limited utility, 16 distinct correct but useless invariants (e.g., those already checked during compilation, such as type checks and constants, Figure 12), and 2 incorrect invariants (e.g., Figure 13). The remaining invariants were repetitions within these categories. This evaluation aligns with ClassInvGen’s validation results, as our validation pipeline also identified 2 incorrect invariants that failed bdd_manager unit tests.
@@ 24 UNREACHABLE();
Figure 9: Z3 developer-written class invariants for bdd_manager class
@@
Figure 10: Correct and useful invariant for bdd_manager class
@@
Figure 11: Ok invariant for bdd_manager class
@@
Figure 12: Correct and useless invariant for bdd_manager class
@@
Figure 13: Incorrect invariant for bdd_manager class
6 Related Work
@@
In this section, we discuss how ClassInvGen relates to previous works on synthesizing program invariants statically, dynamically, and neurally.
6.1 Static approaches
@@
Static techniques, such as interpolation [31] or abstract interpretation [7] perform a symbolic analysis of source code to compute static over-approximations of runtime behavior and represent them as program invariants over suitable domains. These techniques are often used to prove the safety properties of the code. They focus on synthesizing loop invariants and method pre/postconditions, and a few also handle module-level specifications [22]. Given the undecidability of program verification, these techniques scale poorly for real-world programs, especially in the presence of complex data structures and frameworks. In contrast, ClassInvGen can be applied to large codebases to synthesize high-quality class invariants but does not guarantee soundness by construction.
6.2 Dynamic approaches
@@
Dynamic synthesis techniques, such as Daikon [11], DIG [36], SLING [25], and specification mining [1], learn invariants by observing the dynamic behaviors of programs over a set of concrete execution traces. One advantage of these dynamic techniques is that they can be agnostic to the code and generally applicable to different languages. However, these approaches are limited by the templates or patterns over which the invariants can be expressed. DySy [8] employs dynamic symbolic execution to alleviate the problem of fixed templates for bounded executions but resorts to ad-hoc abstraction for loops or recursion. [17] trained models to predict the quality of invariants generated by tools such as Daikon, but they do not generate new invariants. SpecFuzzer [33] generates numerous candidate assertions via fuzzing to construct templates and filters them using Daikon and mutation testing. Finally, Geminus [2] aims at synthesizing sound and complete class invariants representing the set of reachable states, guiding their search using random test cases termed Random Walk.
@@
Unlike these approaches, ClassInvGen can generate a much larger class of invariants, leveraging multimodal inputs, including source code, test cases, comments, and even the naming convention learned from training data, to enhance invariant synthesis. Further, unlike prior dynamic approaches, LLM-based test generation (an active area of research [27, 40, 45]) reduces the need to have a high-quality test suite to obtain the invariants.
@@
For the use case of static verification, learning-based approaches have been used to iteratively improve the quality of the synthesized inductive invariants [14, 15, 37] from dynamic traces. However, these approaches have not been evaluated in real-world programs due to the need for symbolic reasoning.
6.3 Neural approaches
@@
LLM-based invariant synthesis is an emerging area of research with some noteworthy recent contributions. [38] trained a model for zero-shot invariant synthesis, which incurs high training costs and lacks feedback-driven repair. Their approach uses Daikon-generated invariants as both training data and ground truth, which can lead to spurious invariants.
@@
Prior work on nl2postcond [10] prompts LLMs to generate pre- and postconditions of Python and Java benchmarks, illustrating LLMs’ ability to generate high-quality specifications. However, they do not prune incorrect invariants and do not generate class invariants. Combine this work with ClassInvGen would be an interesting direction for future work.
Two very recent neuro–symbolic pipelines extend LLM prompting to other kinds of specifications. [44] combines GPT-4 with bounded-model checking to infer loop invariants: the LLM enumerates candidate predicates, a BMC oracle filters them, and the surviving predicates are re-assembled into provable invariants, yielding a 97 % success rate on 316 numeric-loop benchmarks. [42] (AutoSpec) weaves static slicing and an off-the-shelf program verifier with LLM generation to synthesise function-level contracts; AutoSpec verifies 79 % of heterogeneous benchmarks plus an X.509 parser case study. Both systems rely on static or SMT-based oracles and target scalar loops or procedure specifications, whereas ClassInvGen tackles pointer-rich class/object invariants in idiomatic C++ and validates them chiefly through dynamic test-suite execution plus mutation testing. The different oracle allows our approach to scale to data-structure code bases where precise SMT models are hard to obtain.
@@
For static verification, recent works include the use of LLM for intent formalization from natural language [21], and inferring specifications and inductive program invariants [19, 30]. None of these techniques scale to real-world programs due to the need for complex symbolic reasoning.
7 Limitations